six pears and three apples cost $3.90. two pears and five apples cost $3.30. how much does one pear cost? write one equation for each situation and solve the system of equations.
let p = the cost of one pear
let a = the cost of one apple
a. write a system of equations that can be used to determine the cost of pears and apples.
b. determine the cost of one pear and one apple.

Answer by checkley75(3666)

6P+3A=390
2P+5A=330 NOW MULTIPLY THIS EQUATION BY -3 & ADD.
----------------------------------------------------
-6P-15A=-990
6P+3A=390
-----------------------
-12A=-600
A=-600/-12
A=50 APPLES. NOW SUBSTITUTE 50 FOR A & SOLVE FOR P
6P+3*50=390
6P+150=390
6P=390-150
6P=240
P=240/6
P=40 NUMBER OF PEARS
PROOF
2*40+5*50=330
80+250=330
330=330